Question 1: The following shows data on the average number of customers processed by several bank service units each day. The hourly wage rate is $10, overhead rate is 1.5 times labor cost, and material cost is $5 per customer. b) Suppose a new, more standardized procedure is to be introduced that will enable each employee to process two additional customer per day. Compute the expected labor and multifactor productivity rates for each unit.

Unit Employees Customers per day
A 4 32
B 5 45
C 4 36
D 7 49

a) Compute the labor productivity and the multifactor productivity for each unit. Use an eight-hour day for multifactor productivity

b) Suppose a new, more standardized procedure is to be introduced that will enable each employee to process two additional customer per day. Compute the expected labor and multifactor productivity rates for each unit.

Question 2: Compute the multifactor productivity measure for each of the days shown for production of chocolate bars in Table below. What do the figures suggest? Assume 8-hour day and an hourly rate of $15. Overhead cost is 1.75 times daily labor rate. Material cost is $4 per pound. Day Output Workers Materials (Pounds)

Day Output Workers Material
1 4600 10 50
2 6600 7 80
3 8200 9 120
4 1200 4 55
Answer #1

1. TOAL WAGE = EMPLOYEE * $10 * 8 HOURS

TOTAL MATERIAL = CUSTOMERS * 5

OVERHEAD = 1.5 * LABOR COST


UNIT

TOTAL WAGE

MATERIAL

OVERHEAD

OUTPUT

PARTIAL

MULTIFACTOR

A

320

160

480

32

32 / 320 = 0.1

32 / (320 + 160 + 480 = 0.0333

B

400

225

600

45

45 / 400 = 0.1125

45 / (400 + 225 + 600 = 0.0367

C

320

180

480

36

36 / 320 = 0.1125

36 / (320 + 180 + 480 = 0.0367

D

560

245

840

49

49 / 560 = 0.0875

49 / (560 + 245 + 840 = 0.0298


AFTER THE CHANGE


UNIT

TOTAL WAGE

MATERIAL

OVERHEAD

OUTPUT

PARTIAL

MULTIFACTOR

A

320

200

480

40

40 / 320 = 0.125

40 / (320 + 200 + 480 = 0.04

B

400

275

600

55

55 / 400 = 0.1375

55 / (400 + 275 + 600 = 0.0431

C

320

220

480

44

44 / 320 = 0.1375

44 / (320 + 220 + 480 = 0.0431

D

560

315

840

63

63 / 560 = 0.1125

63 / (560 + 315 + 840 = 0.0367


2.

TOTAL WAGE = WORKERS * 8 HOURS * $15

OVERHEAD = 1.75 * LABOR COST

MATERIALS = PER POUND * $4



DAY

OUTPUT

TOTAL WAGE

MATERIAL

OVERHEAD

MULTIFACTOR

1

4600

1200

200

2100

4600 / (1200 + 200 + 2100) = 1.3143

2

6600

840

320

1470

6600 / (840 + 320 + 1470) = 2.5095

3

8200

1080

480

1890

8200 / (1080 + 480 + 1890) = 2.3768

4

1200

480

220

840

1200 / (480 + 220 + 840) = 0.7792

BASED ON THE MFP, WE CAN SAY THAT THE PRODUCTIVITY DEPENDS HIGHLY OVER MATERIAL COST AND LABOR COST SINCE LABOR HAS THE BIGGEST IMPLICATION OVER THE COST FACTOR, THE MOST OPTIMUM OPTION IS TO UTILIZE A MEDIUM WORKFORCE NUMBER WITH HIGHER MATERIAL BECAUSE OF THE HIGH OVERHEAD COST.

**PLEASE RATE WITH A THUMBS UP :)